Solve this: _+_+_=30 Fill the boxes using (1, 3, 5, 7, 9, 11, 13, 15) U can also repeat the numbers.
Found 3 solutions by macston, Hanssilver, abhik143:
Answer by macston(5194)

No possible answer: The sum of an odd number of odd numbers always results in an odd number.

Answer by Hanssilver(1)

15 + 9 + 6 =30 erse of 9 is 6)

Answer by abhik143(1)

Solutions to Fill the boxes.
I use all numbers one time.
1+3+5+7+9+11+13+15 = 64
64-30 = 34. 34/2 = 17
a+b+c=17 . a = 3, b = 5, c = 9 The boxes are A, B and C.
One solution is A [15+1 -3] +B [13+7 -5] +C [11 - 9] = 30
There are a lot of solutions!
Another solution is : a=1, b= 3, c = 13 A [5+7 - 1] + B [9 - 3] +C [15+11 - 13] = 30
http://puzzle.queryhome.com/6217/solve-this-%E2%97%BB-%E2%97%BB-%E2%97%BB-30?show=7816#a7816